对于事件 AA 和 BB,已知 P(A)=0.55\mathrm{P}(A) = 0.55,P(B)=0.65\mathrm{P}(B) = 0.65 且 P(A∪B)=0.93\mathrm{P}(A \cup B) = 0.93。

  1. (i)

    求 P(A∩B)\mathrm{P}(A \cap B),

  2. (ii)

    求 P(A∪B′)\mathrm{P}(A \cup B')。

  3. (iii)

    判断 AA 和 BB 是否独立。

  4. (iv)

    进一步已知 P(C)=0.3\mathrm{P}(C) = 0.3,且事件 BB 和 CC 是互斥的。

    求 P(A∩C)\mathrm{P}(A \cap C) 的最大值 and 最小值。

先不看答案,让 AI 诊断你卡在哪里?

它不会直接告诉你答案,而是问你卡在哪一步看看 AI 是怎么陪你一步步弄懂这道题的

完整解法

11(i)

1. State the addition rule for probabilities

P(A∪B)=P(A)+P(B)−P(A∩B)\mathrm{P}(A \cup B) = \mathrm{P}(A) + \mathrm{P}(B) - \mathrm{P}(A \cap B)

2. Substitute the given values to find $\mathrm{P}(A \cap B)$

0.93=0.55+0.65−P(A∩B)0.93 = 0.55 + 0.65 - \mathrm{P}(A \cap B)
P(A∩B)=1.20−0.93\mathrm{P}(A \cap B) = 1.20 - 0.93
P(A∩B)=0.27\mathrm{P}(A \cap B) = 0.27

答案

0.270.27

11(ii)

1. Find the probability of the region outside both $A$ and $B$.

P((A∪B)′)=1−P(A∪B)=1−0.93=0.07\mathrm{P}((A \cup B)') = 1 - \mathrm{P}(A \cup B) = 1 - 0.93 = 0.07

2. Use Venn diagram logic to express $A \cup B'$ as the sum of mutually exclusive components and calculate its value.

By visualising a Venn diagram, the region A∪B′A \cup B' consists of the circle AA and the region outside both circles (A∪B)′(A \cup B)'.
Since AA and (A∪B)′(A \cup B)' are mutually exclusive:
P(A∪B′)=P(A)+P((A∪B)′)=0.55+0.07=0.62\mathrm{P}(A \cup B') = \mathrm{P}(A) + \mathrm{P}((A \cup B)') = 0.55 + 0.07 = 0.62

答案

0.620.62

11(iii)

1. Retrieve or calculate $\mathrm{P}(A \cap B)$ from part (i) and calculate $\mathrm{P}(A) \times \mathrm{P}(B)$

From part (i), we have P(A∩B)=P(A)+P(B)−P(A∪B)=0.55+0.65−0.93=0.27\mathrm{P}(A \cap B) = \mathrm{P}(A) + \mathrm{P}(B) - \mathrm{P}(A \cup B) = 0.55 + 0.65 - 0.93 = 0.27.

Calculating the product of the individual probabilities:
P(A)×P(B)=0.55×0.65=0.3575\mathrm{P}(A) \times \mathrm{P}(B) = 0.55 \times 0.65 = 0.3575

2. Compare the values to determine independence

Since 0.27≠0.35750.27 \neq 0.3575, we have P(A∩B)≠P(A)×P(B)\mathrm{P}(A \cap B) \neq \mathrm{P}(A) \times \mathrm{P}(B).
Therefore, events AA and BB are not independent.

答案

Events AA and BB are not independent, since P(A∩B)≠P(A)×P(B)\mathrm{P}(A \cap B) \neq \mathrm{P}(A) \times \mathrm{P}(B).

11(iv)

1. Determine the probabilities of regions available for event $C$.

From part (i), we have P(A∩B)=0.55+0.65−0.93=0.27\mathrm{P}(A \cap B) = 0.55 + 0.65 - 0.93 = 0.27.

Since events BB and CC are mutually exclusive, B∩C=∅B \cap C = \emptyset, which means CC must lie entirely within B′B'.
The region B′B' can be partitioned into two parts: inside AA (A∩B′A \cap B') and outside AA (A′∩B′A' \cap B').

P(A∩B′)=P(A)−P(A∩B)=0.55−0.27=0.28\mathrm{P}(A \cap B') = \mathrm{P}(A) - \mathrm{P}(A \cap B) = 0.55 - 0.27 = 0.28
P(A′∩B′)=1−P(A∪B)=1−0.93=0.07\mathrm{P}(A' \cap B') = 1 - \mathrm{P}(A \cup B) = 1 - 0.93 = 0.07

2. Find the greatest possible value of $\mathrm{P}(A \cap C)$.

For P(A∩C)\mathrm{P}(A \cap C) to be as large as possible, CC should lie inside A∩B′A \cap B' as much as possible.
Since CC cannot overlap with BB, the event A∩CA \cap C must be a subset of A∩B′A \cap B'. Therefore:
P(A∩C)≤P(A∩B′)=0.28\mathrm{P}(A \cap C) \le \mathrm{P}(A \cap B') = 0.28

Since P(C)=0.3\mathrm{P}(C) = 0.3, which is larger than 0.280.28, it is possible for CC to completely cover A∩B′A \cap B'.
Greatest possible value =0.28= 0.28

3. Find the least possible value of $\mathrm{P}(A \cap C)$.

For P(A∩C)\mathrm{P}(A \cap C) to be as small as possible, CC should lie completely outside AA to the maximum extent possible. The only available region for CC outside AA is A′∩B′A' \cap B'.

P(C∩A′)≤P(A′∩B′)=0.07\mathrm{P}(C \cap A') \le \mathrm{P}(A' \cap B') = 0.07

Using the fact that P(C)=P(C∩A)+P(C∩A′)\mathrm{P}(C) = \mathrm{P}(C \cap A) + \mathrm{P}(C \cap A'):
P(A∩C)=P(C)−P(C∩A′)\mathrm{P}(A \cap C) = \mathrm{P}(C) - \mathrm{P}(C \cap A')
P(A∩C)≥0.30−0.07=0.23\mathrm{P}(A \cap C) \ge 0.30 - 0.07 = 0.23

Least possible value =0.23= 0.23

答案

Greatest possible value is 0.280.28
Least possible value is 0.230.23

AI 分析,可能出错

用 AI 批改,圈出写错的那一步,并自动将你的薄弱点存入错题本,形成你的错因库。

同一知识点的其他题

  1. 1已知事件 AA 和 BB 满足 P(B)=13\mathrm{P}(B) = \frac{1}{3},P(A∣B)=25\mathrm{P}(A|B) = \frac{2}{5} 且 P(A∪B)=35\mathrm{P}(A \cup B) = \frac{3}{5}。…
  2. 2凯(Kai)和里奥(Leo)两兄弟是定期训练的双项全能(跑步和骑行)运动员。在每一天中,凯骑行的概率为 34\frac{3}{4},跑步的概率为 35\frac{3}{5},两者都进行的概率为 pp。 平均而言,里奥在一周的 7 天中有…
  3. 3一对同胞 Abbey 和 Betty 还有其他六个朋友一起去参加书展。他们八个人随机排成一队等候一位著名作家的签名。求满足以下条件的概率: Abbey 排在队列第一位且 Betty 排在队列第六位, Abbey 排在队列第一位或 Betty…
  4. 4已知某种特定疾病的感染率约为万分之一(每 10 000 人中约有 1 人感染)。针对该疾病的一项检测准确率为 99.9%,即检测结果与真实情况相符的概率为 99.9%。求在已知某人检测结果呈阳性的条件下,此人确实患有该疾病的概率。 研究人员…
  5. 5对于事件 AA 和 BB,已知 P(A′)=0.3\mathrm{P}(A') = 0.3 且 P(B)=0.55\mathrm{P}(B) = 0.55。 已知事件 AA 和 BB 相互独立,求 P(A∩B)\mathrm{P}(A \cap B)。 已知事件 AA…